# Conformal Temporal Logic Planning using Large Language Models

### v4

@@ III-C LLM-based Planner: How to Accomplish the Assigned Language-based Sub-Task?
To synthesize a finite horizon plan $\tau$ satisfying conditions (a)-(b), we employ pre-trained LLMs. Hereafter, we simply refer to the requirements (a)-(b) as the (language-based) sub-task that the robot should accomplish next. We convert this planning problem into a sequence of $T>0$ multiple-choice question-answering (MCQA) problems for the LLM where $T$ is a hyperparameter. [^1] Essentially, $\tau$ and $T$ refer to $\tau_{n}$ and $T_{n}$ defined in Section II. For ease of presentation, we drop the dependence on the index $n$. The ‘question’ refers to the sub-task along with any actions that the robot has taken up to time $t+k$, $k\in\{0,\dots,T-1\}$ to accomplish it. We denote by $\ell(t+k)$ the (textual) description of the ‘question’. The ‘choices’ refer to the decisions $s(t+k)$ that the robot can make. Given $\ell(t+k)$, the LLM will select $s(t+k)$. This process occurs sequentially for $k\in\{0,\dots,T-1\}$ giving rise to a plan $\tau$. In what follows, we discuss how $\ell(t+k)$ is structured and then we show how it can be used to compute $s(t+k)$.
@@
Figure 2: Example of a prompt constructed in our simulations in Section V. This prompt refers to $t=1$ when the history of actions is empty.
@@
Prompt Construction: In what follows, we discuss how we structure the context text (‘prompt’) that will be given as an input to the LLM; see also Fig. 2. The prompt used in this work consists of the following parts. (A) System description that defines the action space determining all possible actions $a\in{\mathcal{A}}$ that the robot can apply as well as the objective of the LLM. (B) Environment description that describes the locations ${\mathbf{x}}_{e}$ of each semantic object $e$ of interest; (C) Task description that includes the language-based task $\pi_{\text{next}}$ (condition (a)) as well as language-based constraints (if any), modeled by $\Sigma^{\text{unsafe}}$ that the robot should respect until $\pi_{\text{next}}$ is satisfied (condition (b)); (D) History of actions & current environment status that includes the sequence of actions, generated by the LLM, that the robot has executed so far towards accomplishing the assigned task. It also includes the current locations of semantic objects that the robot may have manipulated/moved so far; (E) Response structure describing the desired structure of the LLM output for an example task.
@@
Plan Design & Execution: At iteration $k=0$, part (D) does not include any textual information as a new sub-task has just been announced [line 3, alg. 1]. Given $\ell(t+k)$, the LLM is asked to make a decision $s(t+k)$ among all available ones included in part (A). The LLM selects $s(t+k)$ as follows. Given any $s\in{\mathcal{S}}$, LLMs can provide a score $g(s|\ell(t+k))$; the higher the score, the more likely the decision $s$ is a valid next step to address the language-instruction provided in $\ell(t+k)$. To get $g(s|\ell(t+k))$, we query the LLM over all potential decisions $s\in{\mathcal{S}}$ [50].
@@
Using these scores, a possible approach to select $s(t+k)$ is by simply choosing the decision with the highest score, i.e., $s(t+k)=\arg\max_{s\in\mathcal{S}}g(s|\ell(t+k))$. However, these scores do not represent calibrated confidence. Thus, we calibrate these confidence scores and let the LLM make decisions only when it is certain enough [41]. We formalize this by constructing a set of actions (called, hereafter, prediction set), denoted by ${\mathcal{C}}(\ell(t+k))$, that contain the ground truth action with user-specified confidence [line 7, alg. 1]. Hereafter, we assume that such prediction sets are provided; we defer their construction, using conformal prediction, to Section III-E (see (12)). Given ${\mathcal{C}}(\ell(t+k))$, we select the decision $s(t+k)$ as follows. If $|{\mathcal{C}}(\ell(t+k))|=1$, then we select the decision included in ${\mathcal{C}}(\ell(t+k))$ as it contributes to mission progress with high confidence [line 15, Alg. 1]. [^2] By construction of the prediction sets, this action coincides with $s(t+k)=\arg\max_{s\in\mathcal{S}}g(s|\ell(t+k))$; see Section III-E. As soon as it is selected, the robot physically executes it using its library of low-level controllers [line 23, alg. 1]. If $|{\mathcal{C}}(\ell(t+k))|>1$ [line 8-13, alg.1] or if $|{\mathcal{C}}(\ell(t+k))|=1$ but the robot cannot physically execute $s(t+k)$ the robot seeks assistance to select $s(t+k)$ [line 16-21, Alg.1]. The process of asking for assistance is described in more detail in Section III-D.
@@
Once a decision $s(t+k)$ is made and physically executed, the current time step is updated to $t+k+1$. Then, the prompt $\ell(t+k+1)$ is constructed that will be used to select $s(t+k+1)$. Parts (A)-(C) and (E) in the prompts $\ell(t+k)$ and $\ell(t+k+1)$ are the same. Part (D) in $\ell(t+k+1)$ is augmented by recording the decision $s(t+k)$ as well as incorporating perceptual feedback about the status of containers (if any) that may contain any of the semantic objects. We automatically convert this perceptual feedback into text denoted by $p(t+1)$ (see Fig. 2) [line 24, Alg. 1]. With slight abuse of notation, we denote this prompt update by
$h(t+1)=h(t)+s(t)+p(t+1),$
where the summation means concatenation of text [line 25, Alg.1]. This process is repeated for all $k\in\{0,\dots,T-1\}$ to sequentially select $s(t+k)$. This process generates a plan $\tau$ of length $T$ as in (2) [line 26, alg. 1]. At time $t+T$, part (D) in $\ell(t+T)$ is updated so that it does not include any information since a new sub-task will be announced by the LTL task planner [line 29, alg. 1]. Concatenation of all plans $\tau$ for the sub-tasks generated by the LTL planner gives rise to the plan $\tau_{\phi}$ [line 27, alg. 1] (see Section II).
III-D When to Seek Assistance?
@@
Assume that there exists at least one $k\in\{0,\dots,T-1\}$ so that $|\mathcal{C}(\ell(t+k)))|>1$. In this case, the robot asks for help in order to proceed [line 8-13, Alg. 1]. This assistance request and response occurs as follows. Initially, the robot requests from the LTL task planner an alternative sub-task to make mission progress. To make this help-request, the robot updates the set $\Sigma^{\text{unc}}$, defined in Section III-B, as $\Sigma^{\text{unc}}=\Sigma^{\text{unc}}\cup\{\pi_{\text{next}}\}$. The resulting set $\Sigma^{\text{unc}}$ along with the current mission status are sent to the LTL task planner which then generates a new sub-task [line 9, alg. 1]. If such a sub-task exists, the process discussed in Section III-C repeats [line 11, alg. 1]. If there are no alternative sub-tasks to proceed, then the LTL planner fails to provide assistance. In this case, the robot asks for help from a human operator [line 13, alg. 1]. Specifically, the robot returns to the user the current prompt $\ell(t+k)$ along with the prediction set $\mathcal{C}(\ell(t+k)))$ for which it holds $|\mathcal{C}(\ell(t+k)))|>1$. The user selects the correct decision $s(t+k)$ if it exists in the prediction set or it halts operation. Help from the LTL planner is also requested if the robot cannot physically execute the selected decision $s(t+k)$. This can occur if the robot detects that the initially unknown geometric structure of the environment prevents it from reaching desired destinations/objects, as in [19, 20] [line 16-21, Alg. 1]; see Sec. V-C. In this case, if the LTL planner cannot generate alternative sub-tasks the mission terminates unsuccessfully [line 21, alg. 1].
III-E Constructing Prediction Sets using Conformal Prediction
@@
1:Inputs: Calibration set ${\mathcal{M}}$; New mission scenario; Current prompt ${\ell}_{\text{test}}(t)$; Threshold $\alpha$
@@
In this section, we discuss how the prediction sets ${\mathcal{C}}(\ell(t))$, introduced in Section III-C, are constructed. To construct them, we employ conformal prediction (CP), a statistical tool for uncertainty quantification in black box models [65]. Our analysis builds upon [41, 60]. Construction of these prediction sets requires a distribution ${\mathcal{D}}$ from which LTL-NL missions are sampled. This is important as the application of CP requires the construction of a calibration dataset collecting missions, along with the corresponding ground truth plans, that are independently generated by ${\mathcal{D}}$. Then, the prediction sets for a new/validation mission are ‘correct’ as long as the calibration and the validation missions are i.i.d.; a discussion on relaxing this i.i.d. assumption is provided in Remark IV.4.
[…]
Single-step Plans: We begin by considering LTL-NL formulas $\phi$ that can be satisfied by plans $\tau_{\phi}$ of horizon $H=1$ (see Section II); later we generalize the results for $H\geq 1$. This also means that synthesizing $\tau_{\phi}$ requires the LLM to make a single decision $s$. First, we sample $M$ independent scenarios from ${\mathcal{D}}$. We refer to these scenarios as calibration scenarios. For each calibration scenario $i\in\{1,\dots,M\}$, we construct its equivalent prompt $\ell_{\text{calib}}^{i}$. For each prompt, we (manually) compute the ground truth plan $\tau_{\text{calib}}^{i}=s_{\text{calib}}^{i}(1)$ accomplishing this task. For simplicity, we assume that there exists a unique correct decision $s_{\text{calib}}^{i}$ for each calibration scenario. [^3] This assumption can be relaxed as in [41] yielding prediction sets that contain the correct choice with the highest LLM confidence value (among all other correct choices). Hereafter, we drop the dependence on the robot decisions and prompts on the time step, since we consider single-step plans. This way we construct a calibration dataset =${\mathcal{M}}=\{(\ell_{\text{calib}}^{i},\tau_{\text{calib}}^{i})\}_{i=1}^{M}$.
@@
$P(s_{\text{test}}\in\mathcal{C}(\ell_{\text{test}}))\geq 1-\alpha,$ (4)
@@
where $\alpha\in[0,1]$ is user-specified. To generate $\mathcal{C}(\ell_{\text{test}})$, CP first uses the LLM’s confidence score $g$ (see Section III-C) to compute the set of non-conformity (NC) scores $\{r_{i}=1-g(s_{\text{calib}}^{i}~{}|~{}\ell_{\text{calib}}^{i})\}_{i=1}^{M}$ over the calibration set. The higher the NC score $r_{i}$ is, the worse the performance of the LLM is at the $i$ calibration point. Then CP performs calibration by computing the $\frac{(M+1)(1-\alpha)}{M}$ empirical quantile of $r_{1},\dots,r_{M}$ denoted by $q$. Then, it generates the prediction set
$\mathcal{C}(\ell_{\text{test}})=\{s\in{\mathcal{S}}~{}|~{}g(s|\ell_{\text{test}})>1-q\},$ (5)
@@
that includes all decisions that the predictor is at least $1-q$ confident in. The generated prediction set ensures the $1-\alpha$ marginal coverage guarantee in (4) holds. [^4] To obtain a meaningful prediction set, it should hold $M\geq\big{\lceil}(M+1)(1-\alpha)\big{\rceil}$; otherwise the prediction set will be empty. This coverage guarantee is marginal in the sense that the probability is defined over the randomness of the calibration dataset and the validation scenario. By construction of the prediction sets the decision $s=\arg\max_{s\in\mathcal{S}}g(s|\ell_{\text{test}})$ belongs to $\mathcal{C}(\ell_{\text{test}})$.
@@
Multi-step Plans: Next, we generalize the above result to the case where satisfaction of $\phi$ requires plans $\tau_{\phi}$ with $H\geq 1$ decisions selected from ${\mathcal{S}}$; see Alg. 2. Here we cannot apply directly (4)-(5) to compute individual sets ${\mathcal{C}}(\ell_{\text{test}}(t))$ for the robot, as this violates the i.i.d. assumption required to apply CP. The challenge in this case is that the prompts $\{(\ell_{\text{test}}(t)\}_{t=1}^{H}$ are not independent of each other which violates the i.i.d. assumption required to apply CP. In fact these prompts depend on past robot decisions as well as on the LTL-NL tasks $\phi_{\text{test}}$. To address this challenge, inspired by [41], we (i) lift the data to sequences, and (ii) perform calibration at the sequence level using a carefully designed NC score function.
@@
First, we construct a calibration dataset as follows. We generate $M\geq 1$ scenarios $\xi_{i}$ from ${\mathcal{D}}$. The LTL-NL formula $\phi_{i}$ of each scenario is broken into a sequence of $H_{i}\geq 1$ prompts, defined as: [^5] The distribution ${\mathcal{D}}$ over scenarios induces a distribution over data sequences (6) [41]. These data sequences are equivalent representations of the sampled scenarios augmented with the ground truth decisions.
¯$\bar{\ell}_{\text{calib}}^{i}=[{\ell}_{\text{calib}}^{i}(1),\dots,{\ell}_{\text{calib}}^{i}(H_{i})],$ (6)
@@
where each prompt in the sequence ¯$\bar{\ell}_{\text{calib}}^{i}$ contains a history of ground truth decisions made so far. We define the corresponding sequence of ground truth decisions as:
$\tau_{\phi,\text{calib}}^{i}=s_{\text{calib}}^{i}(1),\dots,s_{\text{calib}}^{i}(H_{i}),$ (7)
@@
This gives rise to the calibration set ={(¯${\mathcal{M}}=\{(\bar{\ell}_{\text{calib}}^{i},\tau_{\phi,\text{calib}}^{i})\}_{i=1}^{M}$. As before, we assume that each context ¯$\bar{\ell}_{\text{calib}}^{i}$ has a unique correct plan $\tau_{\phi,\text{calib}}^{i}$. Next, we use the lowest score $g$ over the time-steps $1,\dots,H_{i}$ as the score for each sequence $i$ in calibration set, i.e.,
$\bar{g}(\tau_{\phi,\text{calib}}^{i}~{}|~{}\bar{\ell}_{\text{calib}}^{i})=\min_{t\in\{1,\dots,H_{i}\}}g(\tau_{\phi,\text{calib}}^{i}(t)~{}|~{}{\ell}_{\text{calib}}^{i}(t)).$ (8)
@@
$\bar{r}_{i}=1-\bar{g}(\tau_{\phi,\text{calib}}^{i}|\bar{\ell}_{\text{calib}}^{i}).$ (9)
@@
¯$\bar{\ell}_{\text{test}}={\ell}_{\text{test}}(1),\dots,{\ell}_{\text{test}}(k),\dots,{\ell}_{\text{test}}(H_{\text{test}}).$ (10)
@@
$P(\tau_{\phi,\text{test}}\in\bar{\mathcal{C}}(\bar{\ell}_{\text{test}}))\geq 1-\alpha,$ (11)
@@
$\bar{\mathcal{C}}(\bar{\ell}_{\text{test}})=\{\tau_{\phi}~{}|~{}\bar{g}(\tau_{\phi}|\bar{\ell}_{\text{test}})>1-\bar{q}\},$ (12)

### v5

@@ 3.3. LLM-based Action Planner: How to Accomplish the Assigned NL-based Sub-Task?
In this section, we present our neural planner, that employs pre-trained LLMs, to synthesize a finite horizon plan $\tau$ satisfying conditions (a)-(b) (see Fig. 3). Hereafter, we simply refer to the requirements (a)-(b) as the (NL-based) sub-task that the robot should accomplish next. Inspired by (Ahn et al., 2022; Ren et al., 2023), we convert this planning problem into a sequence of $T>0$ multiple-choice question-answering (MCQA) problems for the LLM where $T$ is a hyperparameter. [^1] Essentially, $\tau$ and $T$ refer to $\tau_{n}$ and $T_{n}$ defined in Section 2. For ease of presentation, we drop the dependence on $n$. The ‘question’ refers to the sub-task along with any actions that the robot has taken up to time $t+k$, $k\in\{0,\dots,T-1\}$ to accomplish it, where $t$ stands for time step when the sub-task was announced by the symbolic planner. We denote by $\ell(t+k)$ the (textual) description of the ‘question’. The ‘choices’ refer to the decisions $s(t+k)$ that the robot can make. Given $\ell(t+k)$, the LLM will select $s(t+k)$. This process occurs sequentially for $k\in\{0,\dots,T-1\}$ giving rise to a plan $\tau$. Next, we discuss how $\ell(t+k)$ is structured and be used to compute $s(t+k)$.
@@
Figure 4. Example of a prompt constructed in our simulations in Section 5 using GPT 3.5. This prompt refers to $t=1$ when the history of actions is empty.
@@
Prompt Construction: Here we discuss how we structure the context text (‘prompt’) that will be given as an input to the LLM; see also Fig. 4. The prompt in this work consists of the following parts. (A) System description that defines the action space determining all possible actions $a\in{\mathcal{A}}$ that the robot can apply as well as the objective of the LLM. (B) Environment description that describes the initial locations ${\mathbf{x}}_{e}$ of each semantic object $e$ of interest; (C) Task description that includes the language-based task $\pi_{\text{next}}$ (condition (a)) as well as language-based constraints (if any), modeled by $\Sigma_{\text{unsafe}}$ that the robot should respect until $\pi_{\text{next}}$ is satisfied (condition (b)); (D) History of actions & current environment status that includes the sequence of actions, generated by the LLM, that the robot has executed so far towards accomplishing the assigned sub-task. It also includes the current locations of semantic objects that the robot may have manipulated/moved so far; (E) Response structure describing the desired structure of the LLM output for an example task.
@@
Plan Design & Execution: At iteration $k=0$, part (D) does not include any textual information as a new sub-task has just been announced [line 3, Alg. 1]. Given $\ell(t+k)$, the LLM is asked to make a decision $s(t+k)$ among all available ones included in part (A). The LLM selects $s(t+k)$ as follows. Given any $s\in{\mathcal{S}}$, LLMs can provide a logit/confidence score $g(s|\ell(t+k))$; the higher the score, the more likely the decision $s$ is a valid next step to address the language-instruction provided in $\ell(t+k)$. To get $g(s|\ell(t+k))$, we query the LLM over all potential decisions $s\in{\mathcal{S}}$ (Ahn et al., 2022).
@@
Using these scores, a possible approach to select $s(t+k)$ is by simply choosing the decision with the highest score, i.e., $s(t+k)=\arg\max_{s\in\mathcal{S}}g(s|\ell(t+k))$. However, these scores do not represent calibrated confidence. Thus, we calibrate these confidence scores and let the LLM make decisions only when it is certain enough (Ren et al., 2023). We formalize this by constructing a set of actions (called, hereafter, prediction set), denoted by ${\mathcal{C}}(\ell(t+k))$, that contain the ground truth action with user-specified confidence determined by the user-specified mission success rate $1-\alpha$ defined in Section 2 [line 7, Alg. 1]. Hereafter, we assume that such prediction sets are provided; we defer their construction, using conformal prediction, to Section 3.5 (see (11)). Given ${\mathcal{C}}(\ell(t+k))$, we select the decision $s(t+k)$ as follows. If $|{\mathcal{C}}(\ell(t+k))|=1$, then we select the decision included in ${\mathcal{C}}(\ell(t+k))$ as it contributes to mission progress with high confidence [line 20, Alg. 1]. By construction of the prediction sets, this action coincides with $s(t+k)=\arg\max_{s\in\mathcal{S}}g(s|\ell(t+k))$; see Section 3.5. As soon as it is selected, the robot physically executes it using its library of low-level controllers [line 29, Alg. 1]. If $|{\mathcal{C}}(\ell(t+k))|>1$ [line 8-18, Alg.1] or the robot cannot physically execute $s(t+k)$, the robot seeks assistance to select $s(t+k)$ [line 21-27, Alg.1]. The process of asking for assistance is described in more detail in Section 3.4 and illustrated in Fig. 5.
@@
Once a decision $s(t+k)$ is made and executed, the current time step is updated to $t+k+1$. Then, the prompt $\ell(t+k+1)$ is constructed that will be used to select $s(t+k+1)$. Parts (A)-(C) and (E) in the prompts $\ell(t+k)$ and $\ell(t+k+1)$ are the same. Part (D) in $\ell(t+k+1)$ is augmented by recording the decision $s(t+k)$ as well as incorporating perceptual feedback about the status of containers (if any) that may contain any of the semantic objects. We automatically convert this perceptual feedback into text denoted by $p(t+k)$ (see Fig. 4) [line 30, Alg. 1]. With slight abuse of notation, we denote this prompt update by $\ell(t+k+1)=\ell(t+k)+s(t+k)+p(t+k)$ where summation means text concatenation [line 31, Alg.1]. This process is repeated for all $k\in\{0,\dots,T-1\}$ to sequentially select $s(t+k)$. This process generates a plan $\tau$ of length $T$ as in (1) [line 33, Alg. 1].
Figure 5. Graphical overview of the help module. The type of assistance provided depends on the condition that triggered the help mode (case A or B). A ”request new sub-task” message is sent to the symbolic planner (not shown in Fig. 2), while the ”next action” is forwarded to the action execution module in Fig. 3.
At time $t+T$, if the LTL-NL mission has not been accomplished yet, the LTL task planner generates a new sub-task and the above process repeats [line 36, Alg. 1]. The concatenation of all plans $\tau$ for the sub-tasks generated by the LTL planner results in the plan $\tau_{\phi}$ in (2) [line 34, Alg. 1].
3.4. When to Seek Assistance?
@@
Assume that there exists at least one $k\in\{0,\dots,T-1\}$ so that $|\mathcal{C}(\ell(t+k)))|>1$. In this case, the robot asks for help in order to proceed [lines 8-18, Alg. 1]. The help process occurs as follows; see Fig. 5 and Ex. 3.2. Initially, the robot requests help from the human operator [line 9, Alg. 1]. Specifically, the robot presents the current prompt $\ell(t+k)$ to the user along with the prediction set $\mathcal{C}(\ell(t+k)))$. The user selects the correct decision $s(t+k)$ if it exists in the prediction set [line 18, Alg. 1]. If no correct decision is available, the user halts the operation, and the robot will then request from the LTL planner an alternative sub-task to make mission progress [lines 10-16, Alg. 1]. To make this help-request, the set $\Sigma_{\text{unc}}$, defined in Section 3.2, is updated as $\Sigma_{\text{unc}}=\Sigma_{\text{unc}}\cup\{\pi_{\text{next}}\}$ [line 11, Alg. 1]. The resulting set $\Sigma_{\text{unc}}$, along with the current mission status, is sent to the LTL task planner, which then generates a new sub-task [line 12, Alg. 1]. If such a sub-task exists, the process discussed in Section 3.3 repeats [line 14, Alg. 1]. If there are no alternative sub-tasks to proceed, the LTL planner fails to assist. Help from the LTL planner is also requested if the robot cannot physically execute the selected decision $s(t+k)$. This may occur when the robot detects that the initially unknown geometry of the environment prevents it from reaching the desired destinations or objects, as in (Kantaros et al., 2020; Vasilopoulos et al., 2021) [lines 2127, Alg. 1]. In such cases, $\pi_{\text{next}}$ is first added to $\Sigma_{\text{unc}}$ [line 22, Alg. 1], and then the LTL planner is asked to generate an alternative sub-task. If no alternative sub-task can be generated, the mission terminates unsuccessfully [line 27, Alg. 1]. At time step $t+T$, i.e., upon completion of the current sub-task, all sub-tasks previously added to $\Sigma_{\text{unc}}$ due to non-singleton prediction sets not containing the ground truth decision are removed from that set [line 32, Alg. 1]. Instead, sub-tasks added to $\Sigma_{\text{unc}}$ because their execution was physically blocked by the environment remain in the set permanently. This ensures that the symbolic planner will no longer recommend sub-tasks that are physically impossible to complete due to the environment’s geometric structure.
Example 3.2 (Neural Planner & Help Mode).
Consider the LTL-NL formula of Ex. 2.3 and the sub-task, modeled by $\pi_{\text{next}}=\pi_{1}$ and $\Sigma_{\text{unsafe}}=\{\pi_{4}\}$ as generated by the symbolic planner in Ex. 3.1. The neural planner initially generates the action ‘go to Cone 1’ to complete $\pi_{1}$. However, once $\pi_{1}$ is deemed infeasible—due to cone 1 being inaccessible (see Fig. 1)—we request an alternative sub-task from the symbolic planner. In response, the symbolic planner generates $\pi_{\text{next}}=\pi_{3}$, while keeping the unsafe set unchanged, i.e., $\Sigma_{\text{unsafe}}=\{\pi_{4}\}$. The set $\Sigma_{\text{unc}}$, which was initially empty, is updated to $\Sigma_{\text{unc}}(t)=\{\pi_{1}\}$.
3.5. Constructing Prediction Sets
@@
1:Inputs: Calibration set $\mathcal{M}$}; Current prompt ${\ell}_{\text{test}}(t)$; Threshold $\alpha$
@@
In this section, we discuss how the prediction sets ${\mathcal{C}}(\ell(t))$, introduced in Section 3.3, are constructed. To construct them, we employ conformal prediction (CP), a statistical tool for uncertainty quantification in black box models (Angelopoulos et al., 2023). To illustrate the challenges in their construction, we consider two cases: (i) single-step plans, and (ii) multi-step plans. Our analysis builds upon (Ren et al., 2023; Kumar et al., 2023). In what follows, for simplicity, we assume that there exists a unique correct plan for each scenario drawn from $\mathcal{D}$}. This assumption is relaxed in Appendix B.
Single-step Plans: We begin by considering LTL-NL formulas $\phi$ that can be satisfied by plans $\tau_{\phi}$ of horizon $H=1$ (see Section 2); later we generalize the results for $H\geq 1$. This also means that synthesizing $\tau_{\phi}$ requires the LLM to make a single decision $s$. First, we sample $M$ independent scenarios from $\mathcal{D}$}. We refer to these scenarios as calibration scenarios. For each calibration scenario $i\in\{1,\dots,M\}$, we construct its equivalent prompt $\ell_{\text{calib}}^{i}$. For each prompt, we (manually) compute the ground truth plan $\tau_{\text{calib}}^{i}=s_{\text{calib}}^{i}(1)$ accomplishing this task. Hereafter, we drop the dependence on the robot decisions and prompts on the time step, since we consider single-step plans. This way we construct a calibration dataset ${\mathcal{M}}=\{(\ell_{\text{calib}}^{i},\tau_{\text{calib}}^{i})\}_{i=1}^{M}$.
@@
(3) $P(s_{\text{test}}\in\mathcal{C}(\ell_{\text{test}}))\geq 1-\alpha,$
@@
where $\alpha\in(0,1)$ is user-specified. To generate $\mathcal{C}(\ell_{\text{test}})$, CP first uses the LLM’s confidence score $g$ (see Section 3.3) to compute the set of non-conformity (NC) scores $\{r_{i}=1-g(s_{\text{calib}}^{i}~|~\ell_{\text{calib}}^{i})\}_{i=1}^{M}$ over the calibration set. The higher the NC score $r_{i}$ is, the worse the performance of the LLM is at the $i$-th calibration point. Then CP performs calibration by computing the $\frac{(M+1)(1-\alpha)}{M}$ empirical quantile of $r_{1},\dots,r_{M}$ denoted by $q$. Then, it generates the prediction set
(4) $\mathcal{C}(\ell_{\text{test}})=\{s\in{\mathcal{S}}~|~g(s|\ell_{\text{test}})>1-q\},$
@@
that includes all decisions that the predictor is at least 1q1-q confident in. The generated prediction set ensures the $1-\alpha$ marginal coverage guarantee in (3) holds. This coverage guarantee is marginal, defined over the randomness of the calibration data and validation scenario. By construction of the prediction sets the decision $s=\arg\max_{s\in\mathcal{S}}g(s|\ell_{\text{test}})$ belongs to $\mathcal{C}(\ell_{\text{test}})$.
@@
Multi-step Plans: Next, we generalize the above result to the case where satisfaction of $\phi$ requires plans $\tau_{\phi}$ with $H\geq 1$ decisions selected from $\mathcal{S}$}; see Alg. 2. Here we cannot apply directly (3)-(4) to compute individual sets ${\mathcal{C}}(\ell_{\text{test}}(t))$ for the robot, as this violates the i.i.d. assumption required to apply CP. The challenge in this case is that the prompts $\{(\ell_{\text{test}}(t)\}_{t=1}^{H}$ are not independent of each other which violates the i.i.d. assumption required to apply CP. In fact these prompts depend on past robot decisions as well as on the LTL-NL tasks $\phi_{\text{test}}$. To address this challenge, inspired by (Ren et al., 2023), we (i) lift the data to sequences, and (ii) perform calibration at the sequence level using a carefully designed NC score function.
@@
First, we construct a calibration dataset as follows. We generate $M\geq 1$ scenarios $\xi_{i}$ from $\mathcal{D}$}. The LTL-NL formula $\phi_{i}$ of each scenario is broken into a sequence of $H_{i}\geq 1$ prompts, defined as: [^2] The distribution $\mathcal{D}$} over scenarios induces a distribution over data sequences (5) (Ren et al., 2023). These data sequences are equivalent representations of the sampled scenarios augmented with the ground truth decisions.
(5) $\bar{\ell}_{\text{calib}}^{i}=[{\ell}_{\text{calib}}^{i}(1),\dots,{\ell}_{\text{calib}}^{i}(H_{i})],$
@@
where each prompt in the sequence $\bar{\ell}_{\text{calib}}^{i}$ contains a history of ground truth decisions made so far. We define the corresponding sequence of ground truth decisions as:
(6) $\tau_{\phi,\text{calib}}^{i}=s_{\text{calib}}^{i}(1),\dots,s_{\text{calib}}^{i}(H_{i}),$
@@
This gives rise to the calibration set ${\mathcal{M}}=\{(\bar{\ell}_{\text{calib}}^{i},\tau_{\phi,\text{calib}}^{i})\}_{i=1}^{M}$. Next, we use the lowest score $g$ over the time-steps $1,\dots,H_{i}$ as the score for each sequence $i$ in calibration set, i.e.,
(7) $\bar{g}(\tau_{\phi,\text{calib}}^{i}~|~\bar{\ell}_{\text{calib}}^{i})=\min_{t\in\{1,\dots,H_{i}\}}g(\tau_{\phi,\text{calib}}^{i}(t)~|~{\ell}_{\text{calib}}^{i}(t)).$
@@
(8) $\bar{r}_{i}=1-\bar{g}(\tau_{\phi,\text{calib}}^{i}|\bar{\ell}_{\text{calib}}^{i}).$
@@
(9) $\bar{\ell}_{\text{test}}={\ell}_{\text{test}}(1),\dots,{\ell}_{\text{test}}(k),\dots,{\ell}_{\text{test}}(H_{\text{test}}).$
@@
(10) $P(\tau_{\phi,\text{test}}\in\bar{\mathcal{C}}(\bar{\ell}_{\text{test}}))\geq 1-\alpha,$
@@
(11) $\bar{\mathcal{C}}(\bar{\ell}_{\text{test}})=\{\tau_{\phi}~|~\bar{g}(\tau_{\phi}|\bar{\ell}_{\text{test}})>1-\bar{q}\},$